# Deep Research Agents: A Systematic Examination And Roadmap

### v1

@@ Abstract
The rapid progress of Large Language Models (LLMs) has given rise to a new category of autonomous AI systems, referred to as Deep Research (DR) agents. These agents are designed to tackle complex, multi-turn informational research tasks by leveraging a combination of dynamic reasoning, adaptive long-horizon planning, multi-hop information retrieval, iterative tool use, and the generation of structured analytical reports. In this paper, we conduct a detailed analysis of the foundational technologies and architectural components that constitute Deep Research agents. We begin by reviewing information acquisition strategies, contrasting API-based retrieval methods with browser-based exploration. We then examine modular tool-use frameworks, including code execution, multimodal input processing, and the integration of Model Context Protocols (MCPs) to support extensibility and ecosystem development. To systematize existing approaches, we propose a taxonomy that differentiates between static and dynamic workflows, and we classify agent architectures based on planning strategies and agent composition, including single-agent and multi-agent configurations. We also provide a critical evaluation of current benchmarks, highlighting key limitations such as restricted access to external knowledge, sequential execution inefficiencies, and misalignment between evaluation metrics and the practical objectives of DR agents. Finally, we outline open challenges and promising directions for future research. A curated and continuously updated repository of DR agent research is available at: https://github.com/ai-agents-2030/awesome-deep-research-agent.
@@ 1 Introduction
Recent advances in large language models (LLMs) have led to the rapid emergence of sophisticated AI agents capable of autonomous research. Early models such as GPT-3 brown2020languagemodelsfewshotlearners primarily addressed isolated tasks, including question answering and machine translation. Subsequently, integration with external tools enabled models such as WebGPT nakano2021webgpt to navigate the web and synthesize information from diverse sources autonomously. Most recently, a new class of advanced autonomous systems, termed Deep Research (DR) agents, has emerged, exemplified by industry-leading solutions such as OpenAI DR openai2025deepresearch , Gemini DR geminideepresearch , Grok DeepSearch grokdeepresearch , and Perplexity DR perplexitydeepresearch . These deep research agents significantly extend LLMs by incorporating advanced reasoning, dynamic task planning, and adaptive interaction with web resources and analytical tools.
@@
AI agents powered by LLMs, integrating dynamic reasoning, adaptive planning, multi-iteration external data retrieval and tool use, and comprehensive analytical report generation for informational research tasks.
Specifically, DR agents leverage LLMs as their cognitive core, retrieving external knowledge in real-time through web browsers and structured APIs, and dynamically invoking analytical tools via customized toolkits or standardized interfaces such as the Model Context Protocol (MCP). This architecture enables DR agents to autonomously manage complex, end-to-end research workflows by seamlessly integrating reasoning processes with multimodal resources.
Compared with traditional Retrieval-Augmented Generation (RAG) methods singh2025agentic , which primarily enhance factual accuracy but lack sustained reasoning capabilities chen2025improving , and conventional Tool Use (TU) systems qu2025tool that heavily depend on pre-defined workflows wang2025tdag , DR agents offer significantly greater autonomy, continual and deep reasoning abilities, dynamic task planning, and adaptive real-time interaction. These advanced capabilities uniquely position DR agents to handle complex, evolving, and knowledge-intensive research scenarios. A representative example of such a DR agent architecture is illustrated in Figure 1, which demonstrates the complete workflow from user input through optional planning and intent clarification, to iterative tool utilization encompassing offline retrieval (vector and relational databases), online retrieval (APIs and browsers), and extended capabilities including data analytics, coding (etc.), and multimodal generation, ultimately producing comprehensive structured report.
Figure 1: A structural overview of a DR agent in an multi-agent architecture for ease of illustration.
@@
This survey systematically reviews recent advancements in DR agents, providing a comprehensive analysis of core technologies, methodologies, optimization pipelines, and representative implementations. Specifically, the contributions of this survey include:
  • A thorough analysis of representative DR systems, explicitly examining their system architectures, retrieval mechanisms, tool invocation methods, and performance characteristics, alongside optimization and tuning paradigms.
  • A unified classification framework (Figure 4) that systematically categorizes DR systems based on workflow characteristics (static versus dynamic), planning strategies, and agent-based architectures (single-agent versus multi-agent), bridging diverse technical methodologies and current industrial solutions.
  • A systematic review and categorization of existing benchmarks utilized to evaluate DR systems, highlighting how these benchmarks assess critical capabilities, such as retrieval accuracy, reasoning depth, and adaptive tool invocation proficiency.
  • A systematic analysis of critical open challenges and research directions, focusing on expanding retrieval scope beyond traditional methods, enabling asynchronous parallel execution, developing comprehensive multi-modal benchmarks, and optimizing multi-agent architectures for enhanced robustness and efficiency.
@@
This survey methodically explores recent advancements in DR agents, organized as follows: Section 2 provides foundational concepts, examining recent progress in reasoning, retrieval-augmented generation, and agent communication protocols. Section 3 comprehensively analyzes key DR agent components, including search engine integration (Section 3.1), tool invocation strategies (Section 3.2), architectural workflows (Section 3.3), and optimization methodologies (Section 3.4). Section 4 reviews major industrial applications and practical implementations of DR agents by leading organizations. Section 5 surveys benchmarks used for evaluating DR systems, categorizing them into question-answering and task execution scenarios. Section 6 highlights critical challenges and outlines promising directions for future research, focusing on enhancing information acquisition, asynchronous parallel execution, benchmark alignment, and optimizing multi-agent architectures. Finally, Section 7 concludes with a summary and provides insights into the broader implications and opportunities within DR agent research.
@@ 2.1 Advances in Reasoning and Tool Integration
Recent advancements in large reasoning models (LRMs) have greatly enhanced the ability of language models to tackle complex and abstract tasks. These models have shown significant improvements in tasks such as arithmetic, common-sense reasoning, and symbolic problem-solving, largely due to innovations in model architectures and training techniques. One such advancement is Chain-of-Thought (CoT) prompting, introduced by Wei et al. wei2023chainofthoughtpromptingelicitsreasoning , which explicitly guides models to articulate intermediate logical steps, decomposing complex problems into simpler, sequential stages. This has led to notable improvements in both the interpretability and accuracy of LLMs on various reasoning benchmarks. Building upon CoT, subsequent research has introduced methods to further enhance LLM reasoning, particularly in handling lengthy textual contexts. Approaches such as positional interpolation and sparse attention mechanisms bai2024longalign ; wang2024beyond have been proposed to extend the effective context window. Furthermore, specialized benchmarks like LongBench bai2024longbench and LongFinanceQA lin2025facilitating have been developed to rigorously evaluate and improve the performance of these models in extended-context reasoning.
To address reasoning tasks that require real-time or specialized external knowledge, frameworks like Toolformer schick2023toolformerlanguagemodelsteach and MultiTool-CoT inaba2023multitool have been proposed, enabling LLMs to autonomously incorporate external computational resources and APIs directly within reasoning workflows. These approaches effectively enhance performance in tasks dependent on precise numerical calculations and dynamic information retrieval. Maintaining reasoning coherence across multiple conversational turns also poses distinct challenges. Techniques such as Dialogue CoT chae2023dialogue and Structured CoT (SCoT) sultan2024structured explicitly integrate dialogue states and conversational contexts within reasoning chains, significantly improving coherence, context-awareness, and the ability to manage iterative interactions and clarify complex user queries. However, despite substantial improvements, existing reasoning frameworks still encounter critical issues, including hallucinations, static or outdated internal knowledge, and insufficient responsiveness to rapidly changing information needs. These limitations highlight the necessity of integrating external information sources, real-time retrieval mechanisms, and adaptive reasoning strategies, which are core motivations driving recent advances toward more comprehensive and robust reasoning frameworks suitable for DR Agent applications.
@@ 2.2 Advances in Retrieval-Augmented Generation and Agentic Retrieval
Retrieval-augmented Generation (RAG), leveraging external knowledge bases (e.g., webs, APIs), has emerged as an effective strategy to mitigate hallucination problems and enhance the accuracy of web information search fan2024survey ; gao2023retrieval ; singh2025agentic . Early RAG architectures typically involved a static pipeline, where retrievers fetched relevant documents from external sources such as Wikipedia or search engines, and generators (e.g., LLMs) produced answers based solely on these retrieved passages. However, static approaches were limited in handling complex or multi-step queries, motivating recent advances toward iterative and interactive retrieval mechanisms to generate richer and more relevant responses, including FLARE zhang2024enhancing , Self-RAG asai2023self , IAG zhang2023iag , and ToC kim2023tree . In addition, studies izacard2023atlas ; lin2023ra expanded retrieval sources from structured databases (e.g., Wikipedia) to large-scale, diverse web corpora such as the Common Crawl dump preprocessed via the CCNet pipeline fu2022ccnet . Further improvements of RAG include hybrid approaches that combine internal LLM knowledge and external retrievals for better accuracy and coherence aliannejadi2024trec . Recently, Huang et al. huang2025rag proposed RAG-RL, introducing reinforcement learning and curriculum learning techniques, enabling reasoning language models (RLMs) to more effectively identify and utilize relevant contexts.
@@
Despite these advancements in retrieval methods and reasoning-enhanced models, RAG approaches still face limitations in effectively managing complex reasoning workflows and dynamically adapting to varied task requirements. To address these challenges, recent research extends RAG into an agentic paradigm, integrating additional reasoning and decision-making layers atop conventional RAG pipelines singh2025agentic . Agentic RAG approaches leverage iterative retrieval, adaptive querying, and dynamic workflow adjustments, significantly enhancing multi-step reasoning capabilities. For example, RL-based query refinement techniques (e.g., Hsu et al. hsu2024grounding ) improve retrieval for complex queries, while graph-based retrieval (e.g., GeAR shen2024gear ) further enhances the processing of multi-hop queries. Despite these advancements, agentic RAG still faces critical challenges, including balancing computational overhead from dynamic reasoning processes singh2025agentic , aligning agent behaviors with user intentions zerhoudi2024personarag , and ensuring interpretability in adaptive workflows hsu2024grounding ; singh2025agentic . Moreover, even advanced agentic RAG approaches remain constrained by their reliance on pre-existing or periodically updated corpora, limiting their ability to handle real-time, rapidly changing, or long-tail information needs effectively. Addressing this challenge requires integrating external APIs and web browsing capabilities into RAG architectures, motivating recent DR methods aimed at further enhancing retrieval comprehensiveness and adaptability.
@@ 2.3 Model Context Protocol and Agent-to-Agent Policy
Model Context Protocol (MCP) and Agent-to-Agent (A2A) have been proposed to address interoperability challenges in LLM-based agent systems, enabling efficient tool access and effective multi-agent collaboration. MCP: Traditional Tool Use (TU) agents face significant challenges, including inconsistent APIs, high maintenance costs, and redundant development efforts, severely limiting interoperability across systems schick2023toolformerlanguagemodelsteach . To address these issues, Anthropic introduced the MCP, a unified communication layer allowing LLM-based agents to interact securely and consistently with external services and data sources via standardized interfaces. MCP mitigates data silo problems by providing dynamic service discovery and uniform access patterns. A2A: Google’s A2A protocol facilitates decentralized multi-agent collaboration through structured, task-oriented dialogues. Agents from diverse vendors and model architectures can discover peers, delegate responsibilities, and collaboratively manage complex tasks as equal participants google2025a2a . By abstracting agent discovery into Agent Cards, and task coordination into Tasks and Artifacts, A2A supports flexible, incremental, multi-modal workflows, ideally suited to sophisticated collaborative scenarios.
MCP and A2A complement each other by clearly separating responsibilities: MCP serves as a standardized interface for accessing external tools, while A2A orchestrates collaborative agent interactions. Together, they establish a modular and scalable foundation for open, interoperable agent ecosystems, significantly enhancing the practical capabilities of AI systems in tackling complex real-world challenges.
@@ 3 Deep Research: Search Engine, Tool Use, Workflow, Tuning, Non-parametric Continual Learning
Comparison with Coventional RAG-based Approaches. DR agents expand the capabilities of traditional RAG methods by integrating dynamic retrieval, real-time TU, and adaptive reasoning into a unified system. RAG-based approaches typically rely on fixed pipelines, limiting their flexibility in handling complex, multi-step queries or rapidly changing contexts. In contrast, DR agents provide greater autonomy, context-awareness, and accuracy by dynamically engaging with external tools and managing multi-stage research tasks in real time.
In this section, we explore five core components essential for the development and optimization of DR agents: (3.1) search engine integration, which compares API-based interfaces with browser-based exploration to enhance dynamic knowledge acquisition; (3.2) Tool Use capabilities, which investigate the integration of code execution, mathematical computation, file manipulation, and multimodal processing modules within the agent’s inference pipeline; (3.3) workflow architecture, analyzing foundational designs, the balance between multi-agent and single-agent paradigms, memory mechanisms, and auxiliary components that facilitate the orchestration of complex research workflows; (3.4) tuning methodologies, which examine prompt-driven structured generation, LLM-driven prompting, fine-tuning strategies, and reinforcement learning approaches aimed at optimizing agent performance, and (3.5) Non-parametric continual learning, which enables LLM agents to self-evolve by dynamically adapting external tools, memory, and workflows without updating internal model weights, offering scalable optimization for complex tasks.
@@
To enhance reasoning depth and accuracy for handling evolving tasks, DR agents employ search engines (SE) to update their knowledge through interaction with the external environment. In Table 1, we present a comparative overview of SEs, base models, and evaluation benchmarks employed by existing DR agents. The SEs can be broadly categorized into two types:
@@
    API-Based SEs, which interact with structured data sources, such as search-engine APIs or scientific database APIs, enabling efficient retrieval of organized information.
@@
API-based retrieval is a fast, efficient, structured, and scalable method that allows DR agents to access external knowledge sources with relatively less time and computational cost. For instance, Gemini DR leverages multi-source interfaces, most notably the Google Search API and the arXiv API, to perform large-scale retrieval across hundreds to thousands of web pages, thereby significantly expanding its information coverage. Grok DeepSearch grokdeepresearch claims to ensure both the freshness and depth of its knowledge base by maintaining a continuous index via news-outlet feeds, the Wikipedia API, and X’s native interface, and by activating a query-driven agent on demand to generate targeted sub-queries and fetch relevant pages in real time. AgentLaboratory schmidgall2025agent uses the arXiv API to extract paper metadata and abstracts for automated literature reviews. AI Scientist lu2024aiscientistfullyautomated issues requests to the Semantic Scholar API to validate the novelty and citation relationships of model-generated research ideas, and CoSearchAgent gong2024cosearchagent integrates SerpApi to deliver Slack-based, real-time search responses. DeepRetrieval jiang2025deepretrieval , operating within a reinforcement-learning framework, optimizes queries against the PubMed and ClinicalTrials.gov APIs to maximize recall on biomedical tasks, and Search-o1 li2025search combines the Bing Search API with the Jina Reader API to dynamically extract and refine passages for downstream reasoning. Whilst these API-driven methods excel at structured, high-throughput data acquisition, they generally struggle when faced with deeply nested client-side JavaScript-rendered content, interactive components, or authentication barriers, thereby motivating the development of browser-based search mechanisms capable of comprehensively extracting and analyzing dynamic or unstructured information.

### v2

@@ Abstract
The rapid progress of Large Language Models (LLMs) has given rise to a new category of autonomous AI systems, referred to as Deep Research (DR) agents. These agents are designed to tackle complex, multi-turn informational research tasks by leveraging a combination of dynamic reasoning, adaptive long-horizon planning, multi-hop information retrieval, iterative tool use, and the generation of structured analytical reports. In this paper, we conduct a detailed analysis of the foundational technologies and architectural components that constitute Deep Research agents. We begin by reviewing information acquisition strategies, contrasting API-based retrieval methods with browser-based exploration. We then examine modular tool-use frameworks, including code execution, multimodal input processing, and the integration of Model Context Protocols (MCPs) to support extensibility and ecosystem development. To systematise existing approaches, we propose a taxonomy that differentiates between static and dynamic workflows, and we classify agent architectures based on planning strategies and agent composition, including single-agent and multi-agent configurations. We also provide a critical evaluation of current benchmarks, highlighting key limitations such as restricted access to external knowledge, sequential execution inefficiencies, and misalignment between evaluation metrics and the practical objectives of DR agents. Finally, we outline open challenges and promising directions for future research. A curated and continuously updated repository of DR agent research is available at: https://github.com/ai-agents-2030/awesome-deep-research-agent.
@@ 1 Introduction
Recent advances in large language models (LLMs) have led to the rapid emergence of sophisticated AI agents capable of autonomous research. Early models such as GPT-3 (brown2020languagemodelsfewshotlearners, ) primarily addressed isolated tasks, including question answering and machine translation. Subsequently, integration with external tools enabled models such as WebGPT (nakano2021webgpt, ) to navigate the web and synthesise information from diverse sources autonomously. Most recently, a new class of advanced autonomous systems, termed Deep Research (DR) agents, has emerged, exemplified by industry-leading solutions such as OpenAI DR (openai2025deepresearch, ), Gemini DR (geminideepresearch, ), Grok DeepSearch (grokdeepresearch, ), and Perplexity DR (perplexitydeepresearch, ). These deep research agents significantly extend LLMs by incorporating advanced reasoning, dynamic task planning, and adaptive interaction with web resources and analytical tools.
@@
AI agents powered by LLMs, integrating dynamic reasoning, adaptive planning, and iterative tool use to acquire, aggregate, and analyse external information, culminating in comprehensive outputs for accomplishing open-ended informational research tasks.
Specifically, DR agents leverage LLMs as their cognitive core, retrieving external knowledge in real-time through web browsers and structured APIs, and dynamically invoking analytical tools via customised toolkits or standardised interfaces such as the Model Context Protocol (MCP). This architecture enables DR agents to autonomously manage complex, end-to-end research workflows by seamlessly integrating reasoning processes with multimodal resources.
Compared with traditional Retrieval-Augmented Generation (RAG) methods (singh2025agentic, ), which primarily enhance factual accuracy but lack sustained reasoning capabilities (chen2025improving, ), and conventional Tool Use (TU) systems (qu2025tool, ) that heavily depend on pre-defined workflows (wang2025tdag, ), DR agents offer significantly greater autonomy, continual and deep reasoning abilities, dynamic task planning, and adaptive real-time interaction. These advanced capabilities uniquely position DR agents to handle complex, evolving, and knowledge-intensive research scenarios. A representative example of such a DR agent architecture is illustrated in Figure 1, which demonstrates the complete workflow from user input through optional planning and intent clarification, to iterative tool utilization encompassing offline retrieval (vector and relational databases), online retrieval (APIs and browsers), and extended capabilities including data analytics, coding (etc.), and multimodal generation, ultimately producing comprehensive structured report.
Figure 1: A structural overview of a DR agent in a multi-agent architecture for ease of illustration.
@@
This survey systematically reviews recent advancements in DR agents, providing a comprehensive analysis of core technologies, methodologies, optimisation pipelines, and representative implementations. Specifically, the contributions of this survey include:
  • A thorough analysis of representative DR systems, explicitly examining their system architectures, retrieval mechanisms, tool invocation methods, and performance characteristics, alongside optimisation and tuning paradigms.
  • A unified classification framework (Figure 4) that systematically categorises DR systems based on workflow characteristics (static versus dynamic), planning strategies, and agent-based architectures (single-agent versus multi-agent), bridging diverse technical methodologies and current industrial solutions.
  • A systematic review and categorisation of existing benchmarks utilised to evaluate DR systems, highlighting how these benchmarks assess critical capabilities, such as retrieval accuracy, reasoning depth, and adaptive tool invocation proficiency.
  • A systematic analysis of critical open challenges and research directions, focusing on expanding retrieval scope beyond traditional methods, enabling asynchronous parallel execution, developing comprehensive multi-modal benchmarks, and optimising multi-agent architectures for enhanced robustness and efficiency.
@@
This survey methodically explores recent advancements in DR agents, organised as follows: Section 2 provides foundational concepts, examining recent progress in reasoning, retrieval-augmented generation, and agent communication protocols. Section 3 comprehensively analyses key DR agent components, including search engine integration (Section 3.1), tool invocation strategies (Section 3.2), architectural workflows (Section 3.3), and optimisation methodologies (Section 3.4). Section 4 reviews major industrial applications and practical implementations of DR agents by leading organisations. Section 5 surveys benchmarks used for evaluating DR systems, categorising them into question-answering and task execution scenarios. Section 6 highlights critical challenges and outlines promising directions for future research, focusing on enhancing information acquisition, asynchronous parallel execution, benchmark alignment, and optimising multi-agent architectures. Finally, Section 7 concludes with a summary and provides insights into the broader implications and opportunities within DR agent research.
@@ 2.1 Advances in Reasoning and Tool Integration
Recent advancements in large reasoning models (LRMs) have greatly enhanced the ability of language models to tackle complex and abstract tasks. These models have shown significant improvements in tasks such as arithmetic, common-sense reasoning, and symbolic problem-solving, largely due to innovations in model architectures and training techniques. One such advancement is Chain-of-Thought (CoT) prompting, introduced by Wei et al. (wei2023chainofthoughtpromptingelicitsreasoning, ), which explicitly guides models to articulate intermediate logical steps, decomposing complex problems into simpler, sequential stages. This has led to notable improvements in both the interpretability and accuracy of LLMs on various reasoning benchmarks. Building upon CoT, subsequent research has introduced methods to further enhance LLM reasoning, particularly in handling lengthy textual contexts. Approaches such as positional interpolation and sparse attention mechanisms (bai2024longalign, ; wang2024beyond, ) have been proposed to extend the effective context window. Furthermore, specialised benchmarks like LongBench (bai2024longbench, ) and LongFinanceQA (lin2025facilitating, ) have been developed to rigorously evaluate and improve the performance of these models in extended-context reasoning.
To address reasoning tasks that require real-time or specialised external knowledge, frameworks like Toolformer (schick2023toolformerlanguagemodelsteach, ) and MultiTool-CoT (inaba2023multitool, ) have been proposed, enabling LLMs to autonomously incorporate external computational resources and APIs directly within reasoning workflows. These approaches effectively enhance performance in tasks dependent on precise numerical calculations and dynamic information retrieval. Maintaining reasoning coherence across multiple conversational turns also poses distinct challenges. Techniques such as Dialogue CoT (chae2023dialogue, ) and Structured CoT (SCoT) (sultan2024structured, ) explicitly integrate dialogue states and conversational contexts within reasoning chains, significantly improving coherence, context-awareness, and the ability to manage iterative interactions and clarify complex user queries. However, despite substantial improvements, existing reasoning frameworks still encounter critical issues, including hallucinations, static or outdated internal knowledge, and insufficient responsiveness to rapidly changing information needs. These limitations highlight the necessity of integrating external information sources, real-time retrieval mechanisms, and adaptive reasoning strategies, which are core motivations driving recent advances toward more comprehensive and robust reasoning frameworks suitable for DR Agent applications.
@@ 2.2 Advances in Retrieval-Augmented Generation and Agentic Retrieval
Retrieval-augmented Generation (RAG), leveraging external knowledge bases (e.g., webs, APIs), has emerged as an effective strategy to mitigate hallucination problems and enhance the accuracy of web information search (fan2024survey, ; gao2023retrieval, ; singh2025agentic, ). Early RAG architectures typically involved a static pipeline, where retrievers fetched relevant documents from external sources such as Wikipedia or search engines, and generators (e.g., LLMs) produced answers based solely on these retrieved passages. However, static approaches were limited in handling complex or multi-step queries, motivating recent advances toward iterative and interactive retrieval mechanisms to generate richer and more relevant responses, including FLARE (zhang2024enhancing, ), Self-RAG (asai2023self, ), IAG (zhang2023iag, ), and ToC (kim2023tree, ). In addition, studies (izacard2023atlas, ; lin2023ra, ) expanded retrieval sources from structured databases (e.g., Wikipedia) to large-scale, diverse web corpora such as the Common Crawl dump preprocessed via the CCNet pipeline (fu2022ccnet, ). Further improvements of RAG include hybrid approaches that combine internal LLM knowledge and external retrievals for better accuracy and coherence (aliannejadi2024trec, ). Recently, Huang et al. (huang2025rag, ) proposed RAG-RL, introducing reinforcement learning and curriculum learning techniques, enabling reasoning language models (RLMs) to more effectively identify and utilise relevant contexts.
@@
Despite these advancements in retrieval methods and reasoning-enhanced models, RAG approaches still face limitations in effectively managing complex reasoning workflows and dynamically adapting to varied task requirements. To address these challenges, recent research extends RAG into an agentic paradigm, integrating additional reasoning and decision-making layers atop conventional RAG pipelines (singh2025agentic, ). Agentic RAG approaches leverage iterative retrieval, adaptive querying, and dynamic workflow adjustments, significantly enhancing multi-step reasoning capabilities. For example, RL-based query refinement techniques (e.g., Hsu et al. (hsu2024grounding, )) improve retrieval for complex queries, while graph-based retrieval (e.g., GeAR (shen2024gear, )) further enhances the processing of multi-hop queries. Despite these advancements, agentic RAG still faces critical challenges, including balancing computational overhead from dynamic reasoning processes (singh2025agentic, ), aligning agent behaviours with user intentions (zerhoudi2024personarag, ), and ensuring interpretability in adaptive workflows (hsu2024grounding, ; singh2025agentic, ). Moreover, even advanced agentic RAG approaches remain constrained by their reliance on pre-existing or periodically updated corpora, limiting their ability to handle real-time, rapidly changing, or long-tail information needs effectively. Addressing this challenge requires integrating external APIs and web browsing capabilities into RAG architectures, motivating recent DR methods aimed at further enhancing retrieval comprehensiveness and adaptability.
@@ 2.3 Model Context Protocol and Agent-to-Agent Policy
Model Context Protocol (MCP) and Agent-to-Agent (A2A) have been proposed to address interoperability challenges in LLM-based agent systems, enabling efficient tool access and effective multi-agent collaboration. MCP: Traditional Tool Use (TU) agents face significant challenges, including inconsistent APIs, high maintenance costs, and redundant development efforts, severely limiting interoperability across systems (schick2023toolformerlanguagemodelsteach, ). To address these issues, Anthropic introduced the MCP, a unified communication layer allowing LLM-based agents to interact securely and consistently with external services and data sources via standardised interfaces. MCP mitigates data silo problems by providing dynamic service discovery and uniform access patterns. A2A: Google’s A2A protocol facilitates decentralised multi-agent collaboration through structured, task-oriented dialogues. Agents from diverse vendors and model architectures can discover peers, delegate responsibilities, and collaboratively manage complex tasks as equal participants (google2025a2a, ). By abstracting agent discovery into Agent Cards and task coordination into Tasks and Artefacts, A2A supports flexible, incremental, multi-modal workflows, ideally suited to sophisticated collaborative scenarios.
MCP and A2A complement each other by clearly separating responsibilities: MCP serves as a standardised interface for accessing external tools, while A2A orchestrates collaborative agent interactions. Together, they establish a modular and scalable foundation for open, interoperable agent ecosystems, significantly enhancing the practical capabilities of AI systems in tackling complex real-world challenges.
@@ 3 Deep Research: Search Engine, Tool Use, Workflow, Tuning, Non-parametric Continual Learning
Comparison with Conventional RAG-based Approaches. DR agents expand the capabilities of traditional RAG methods by integrating dynamic retrieval, real-time TU, and adaptive reasoning into a unified system. RAG-based approaches typically rely on fixed pipelines, limiting their flexibility in handling complex, multi-step queries or rapidly changing contexts. In contrast, DR agents provide greater autonomy, context-awareness, and accuracy by dynamically engaging with external tools and managing multi-stage research tasks in real time.
In this section, we explore five core components essential for the development and optimization of DR agents: (3.1) search engine integration, which compares API-based interfaces with browser-based exploration to enhance dynamic knowledge acquisition; (3.2) Tool Use capabilities, which investigate the integration of code execution, mathematical computation, file manipulation, and multimodal processing modules within the agent’s inference pipeline; (3.3) workflow architecture, analysing foundational designs, the balance between multi-agent and single-agent paradigms, memory mechanisms, and auxiliary components that facilitate the orchestration of complex research workflows; (3.4) tuning methodologies, which examine prompt-driven structured generation, LLM-driven prompting, fine-tuning strategies, and reinforcement learning approaches aimed at optimizing agent performance, and (3.5) Non-parametric continual learning, which enables LLM agents to self-evolve by dynamically adapting external tools, memory, and workflows without updating internal model weights, offering scalable optimization for complex tasks.
@@
To enhance reasoning depth and accuracy for handling evolving tasks, DR agents employ search engines (SE) to update their knowledge through interaction with the external environment. In Table LABEL:tab:search_engine, we present a comparative overview of SEs, base models, and evaluation benchmarks employed by existing DR agents. The SEs can be broadly categorised into two types:
@@
    API-Based SEs, which interact with structured data sources, such as search-engine APIs or scientific database APIs, enabling efficient retrieval of organised information.
@@
Towards an AI co-scientist gottweis2025towards GPQA Gemini 2.0 Feb-2025
Nouswise nouswise2025 Mar-2025
@@
Tool-Star dong2025tool WebWalker, HotpotQA, 2WikiMultiHopQA Qwen-2.5 May-2025
WebDancer wu2025webdancer WebWalkerQA Qwen-2.5, QwQ-32B, DeepSeek-R1, GPT-4o May-2025
@@
O-agents zhu2025oagents GPT-4o, GPT-4.1, Claude-3.7-Sonnet, DeepSeek-R1, Gemini-2.5 Jun-2025
Kimi-Researcher kimi2025researcher SimpleQA Kimi k1.5/k2 Jun-2025
WebSailor li2025websailor SimpleQA Qwen-2.5 Jul-2025
Agent-KB tang2025agent SWE-bench GPT-4o, GPT-4.1, Claude-3.7-Sonnet, o3-mini, Qwen-3, DeepSeek-R1 Jul-2025
WebShaper tao2025webshaper WebWalkerQA Qwen-2.5, QwQ-32B Jul-2025
Deep Researcher with Test-Time Diffusion han2025deep Gemini-2.5-Pro Jul-2025
ChatGPT-Agent openai2025chatgptagent Jul-2025
AWorld aworld2025 HotpotQA Gemini-2.5-Pro, GPT-4o Jul-2025
Cognitive Kernel-Pro wan2025cognitive AgentWebQA, WebWalkerQA, Multi-hop URLQA, DocBench, TableBench Claude-3.7-Sonnet, CK-Pro-8B Aug-2025
WebWatcher geng2025webwatcher Browsercom-VL, LiveVQA, MMSearch Qwen-2.5-VL-32B Aug-2025
WideSearch wong2025widesearch WideSearch DeepSeek-R1, Doubao-Seed-1.6, Claude Sonnet 4, Gemini-2.5-Pro Aug-2025
MiroRL 2025mirorl Qwen3-14B Aug-2025
@@
API-based retrieval is a fast, efficient, and scalable way for deep-research (DR) agents to access external knowledge with relatively low latency and computational overhead. For example, Gemini DR (geminideepresearch, ) coordinates multiple interfaces, most notably the Google Search and arXiv APIs, to conduct large-scale retrieval over hundreds to thousands of web pages, substantially broadening coverage. Grok DeepSearch (grokdeepresearch, ) maintains a continuously updated index via news-outlet feeds, the Wikipedia API, and X’s native interface, and, on demand, dispatches a query-driven agent to decompose questions into targeted subqueries and fetch relevant pages in real time. perplexitydeepresearch first crawls hundreds of sources and then aggregates them to produce a final report. Cognitive Kernel-Pro (wan2025cognitive, ) leverages the free DuckDuckGo search interface to enable a fully open-source, low-cost DR pipeline. Agentic Reasoning, ReSearch, R1-Search and SWIRL (wu2025agentic, ; chen2025learning, ; song2025r1, ; goldie2025synthetic, ) explicitly teach models when to search, what to search for, and how to incorporate retrieved evidence into the reasoning process. PANGU DeepDiver (shi2025pangu, ) uses reinforcement learning to adapt search intensity to task difficulty. Agent Laboratory (schmidgall2025agent, ) calls the arXiv API to extract paper metadata and abstracts for automated literature reviews, while AI Scientist (lu2024aiscientistfullyautomated, ) queries the Semantic Scholar API to validate novelty and citation relations among model-generated ideas. CoSearch-Agent (gong2024cosearchagent, ) integrates SerpApi to deliver Slack-based, real-time search. DeepRetrieval (jiang2025deepretrieval, ) operates within a reinforcement-learning framework to optimise queries against the PubMed and ClinicalTrials.gov APIs for high-recall biomedical retrieval, and Search-o1 (li2025search, ) combines the Bing Search API with the Jina Reader API to dynamically extract and refine passages for downstream reasoning. While these API-driven approaches excel at structured, high-throughput acquisition, they can struggle with deeply nested, client-side JavaScript–rendered content, interactive components, or authentication barriers, motivating complementary browser-based mechanisms capable of comprehensively extracting and analysing dynamic or unstructured information.